Check my work de up of Main, Capitel $392000, and Frist he Struter Portniership has total partners Captat $168 000.The partners share net income and loss in a ratio of 79% to Main and 21% to Frist on November 1 Adson is shined to the partnership and given 20% inserest in equity and a 20% share in any income and loss. equny of $560,000, which is ma Prepare partial entries to record ee odmission of Adison for a 20% rnerest in the equity and a 20% share in any income Recoed the admission of Adison with an t of $140,000 fore 20% nterest in the equity and อ 20% share in investment (2í nocird trui admission or Adson with en mesinent of $175,000 for a 20% interest r, the equty and a 20% share in ny income and loss uecoro mi, kims sion of Adison with an investment of $10,000 for a 20% interest in the equity and e 20% snare in C Prev
Answer #1
S. No. Particular Debit Credit
1. cash $140,000
Adison capital $140,000
2. Cash $175,000
Adison capital $147,000
Main capital $22,120
First capital $5,880
3. cash $110,000
Main capital $18,960
First capital $5,040
Adison capital $134,000

Working Notes

1.

($560,000+$140,000)*20% = $140,000

2.

($560,000+$175,000)*20% = $147,000

($175,000-$147,000) = $28,000

Main capital = $28,000*79% = $22,120

First capital = $28,000*21% = $5,880

3.

($560,000+$110,000)*20% = $134,000

($110,000-$134,000) = $(24,000)

Main capital = $24,000*79% = $18,960

First capital = $24,000*21% = $5,040